Testing [8, 13, 2, 3, 11, 6, 10, 12, 1, 7, 14, 4, 9, 5]

References

[1] K. Åström, P. Albertos, M. Blanke, A. Isidori, W. Schaufelberger, and R. Sanz, editors. Control of Complex Systems. Springer-Verlag, London, Great Britain, 2001.

[2] M. Brown and C. Harris. Neurofuzzy Adaptive Modelling and Control. Prentice Hall International Series in Systems and Control Engineering, M. J. Grimble, Series Editor. Prentice Hall International (UK) Limited, Hemel Hempstead, Hertfordshire, HP2 7EZ, 1994.

[3] C. J. Harris, C. G. Moore, and M. Brown. Intelligent Control: Aspects of Fuzzy Logic and Neural Nets. World Scientific Series in Robotics and Automated Systems—Vol. 6. World Scientific, Singapore, 1993.

[4] E. J. Klein and W. F. Ramirez. State controllability and optimal regulator control of time-delayed systems. Int. J. Control, 74(3):281–289, 15 February 2001.

[5] F. L. Lewis and K. Liu. Towards a paradigm for fuzzy logic control. Automatica, 32(2):167–181, 1996.

[6] H. Li and M. Gupta, editors. Fuzzy Logic and Intelligent Systems. International Series in Intelligent Technologies. Kluwer Academic Publishers, Boston, 1995.

[7] A. G. J. MacFarlane. Information, knowledge, and machines. In C. W. de Silva, editor, Intelligent Machines: Myths and Realities, pages 43–76, Boca Raton, Florida 33431, 2000. CRC Press. Chapter 2.

[8] G. W. Ng. Application of Neural Networks to Adaptive Control of Nonlinear Systems. UMIST Control Systems Centre Series. RSP Ltd., Taunton, Somerset, England TA1 1HD, 1997.

[9] Y. Osana and M. Hagiwara. Successive learning in hetero-associative memory using chaotic neural networks. Int. J. Neural Systems, 9(4):285–299, August 1999.

[10] D. Ruan, editor. Intelligent Hybrid Systems: Fuzzy Logic, Neural Networks, and Genetic Algorithms. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1997.

[11] D. Schröder, editor. Intelligent Observer and Control Design for Nonlinear Systems. Springer-Verlag, Berlin, Germany, 2000.

[12] S. G. Tzafestas, editor. Computational Intelligence in Systems and Control Design and Applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1999.

[13] D. A. White and D. A. Sofage, editors. Handbook of Intelligent Control: Neural, Fuzzy, and Adaptive Approaches. Van Nostrand Reinhold, New York, 1992.

[14] S. Yashiki and N. Matsumoto. Stabilization and optimal regulator problem for timedelay systems based on the 2D Riccati matrix equation. Electronics and Communications in Japan, Part 3, 81(1):1–12, January 1998.